Required:

 1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter.

 2. What is the accounts receivable balance on June 30th?


Exercise 8-1 Schedule of Expected Cash Collections [LO8-2] 


Silver Company makes a product that is very popular as a Mother's Day gift. Thus, peak sales occur in May of each year, as shown in the company's sales budget for the second quarter given below: 


AprilMayJuneTotal
Budgeted sales (all on account)$300,000$500,000$200,000$1,000,000


From past experience, the company has learned that 20% of a month's sales are collected in the month of sale, another 70% are collected in the month following sale, and the remaining 10% are collected in the second month following sale. Bad debts are negligible and can be ignored. February sales totaled $230,000, and March sales totaled $260,000.Required: 1. Prepare a schedule of expected cash collections from sales, by month and in total, for the second quarter. 2. WhExercise 8-1 Schedule of Expected Cash Collections [LO8-2] Silver Company makes a product that is very popular as a Mothers
Answer #1

Requirement 1

Schedule of Expected Cash Collections
April May June Total
February Sales $ 23,000 $ 23,000
March Sales $ 182,000 $ 26,000 $ 208,000
April Sales $ 60,000 $ 210,000 $ 30,000 $ 300,000
May Sales $ 100,000 $ 350,000 $ 450,000
June Sales $ 40,000 $ 40,000
Total Cash Collection $ 265,000 $ 336,000 $ 420,000 $ 1,021,000

Requirement 2

Total Receivables at June 30 $ 210,000

Working

Silver company
Accounts receivables at june 30
May sales $ 50,000
June sales $ 160,000
Total Accounts receivables at june 30 $ 210,000

.

Accounts Receivables Balance
Balance of may sales (10% of sale) $ 50,000
Balance of June Sale (80% of sale) $ 160,000
Total Receivables at June 30 $ 210,000

.

Schedule of Expected Cash Collections
April May June
February Sales =230000*10%
March Sales =260000*70% =260000*10%
April Sales =300000*20% =300000*70% =300000*10%
May Sales =500000*20% =500000*70%
June Sales =200000*20%
Answer #2
Managerial
source: Managerial accouting
answered by: Faiza